OpenStudy (anonymous):

2x-5y=3 3x-5y=7 solve using elimination method

OpenStudy (callisto):

2x-5y=3 -(1) 3x-5y=7 -(2) (2) - (1): (3x-5y) - (2x-5y) = 7-3 x = 4 Can you solve y?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

first equation is 2x-5y=3 and second one is 3x-5y=7 now we need to eliminate either x or y. If you can see both are -5y so it is easier to eliminate y Multiply one of the equations with (-) 2x-5y=3 -(3x-5y=7) -3x+5y=-7 add both equations 2x-5y=3 + -3x+5y=-7 -x=-4 x=4 plug this in either one of the equations 2x-5y=3 2(4)-5y=3 -5y=-5y y=1 answer :D

OpenStudy (anonymous):

how about 2x-5y=18 -6x+5y=14

OpenStudy (callisto):

This time, add the 2 equations together :) Can you try?

OpenStudy (anonymous):

2 -5 3 3 -5 7 2 -5 3 0 5 5 now back substitution 5y=5 : y=1 x=4 method called guasse's method got?

OpenStudy (anonymous):

x= -4

OpenStudy (callisto):

2x-5y=18 -(1) -6x+5y=14 -(2) (1)+(2) (2x-5y) + (-6x+5y) = 18 +14 -4x = 32 x=?

OpenStudy (anonymous):

that was supposed to be -6x+5y=-14

OpenStudy (anonymous):

my bad

OpenStudy (callisto):

2x-5y=18 -(1) -6x+5y=-14 -(2) (1)+(2) (2x-5y) + (-6x+5y) = 18 +(-14) -4x = 4 x=?

OpenStudy (anonymous):

so when adding i get -4x=4... divid 4 by -4... x =-4

OpenStudy (callisto):

4 /( -4) =?

OpenStudy (anonymous):

o im retarded x=-1

OpenStudy (callisto):

Yes :) and can you do y?

OpenStudy (anonymous):

y=-4

OpenStudy (callisto):

I got the same answer as yours :)